kacchanforlife:

PLSS HELP ASAP!!! 1) What is the equation in point slope form of the line that passes through the point (1 , 8) and (2 , 10)? y+8=2(x+1) y-8=2(x-1) 2x-y=7 y=2x-7 2) If you have points (1,1) and (5,9) on a graph, what is the point slope form? y-1=2(x-1) y+1=2(x+1) 3) If you have points (3,6) and (2,1) on a graph, what is the equation of the line in standard form? 7x+y=13 y-8=7(x-3) 7x-y=13 4) If you have points (1,4) and (-1,2) on a graph, what is the equation of the line in slope intercept form? 2x-y=-3 y=x+3 (more below!)

kacchanforlife:

On the second one, the points are (2,15) and (0,5)....

supie:

@narad this should be light work for you

kacchanforlife:

r they online rn??

supie:

yeah but idk if he’s active

kacchanforlife:

ok ty!

supie:

if not I could help you with some

doggofroyo:

@supie

supie:

okay so for question 1, the equations are written in point slope form, y-y1=x(x-x1) so i would just find the equation that matches y1 and x1 with (1,8) and (2,10)

supie:

ykwim or no?

kacchanforlife:

ya i do!

supie:

nvm i read the question wrong

kacchanforlife:

oh no!

supie:

@aiden3j

Aiden3j:

okay Kacchan do you now how to find a slope?

kacchanforlife:

urm kinda kinda not (sorry its been a while since we've worked on it)

Aiden3j:

ok so the first three steps Locate two points on the line whose coordinates are integers. Starting with the point on the left, sketch a right triangle, going from the first point to the second point. Count the rise and the run on the legs of the triangle. Take the ratio of rise to run to find the slope. m=riserun.

kacchanforlife:

on the 1st one??

kacchanforlife:

m=riserun=ΔyΔx m=y2−y1x2−x1 m=10−82−1 m=21 m = 2 is this right @aiden3j ??

lovemelee:

The point slope form of this equation would be y - 8 = 2(x - 1)

lovemelee:

the slope formula. m = (y2 - y1)/(x2 - x1) m = (10 - 8)/(2 - 1) m = 2/1 m = 2

Aiden3j:

thanks @lovemelee i have to go for a bit

lovemelee:

oh?

kacchanforlife:

hold on imma close this and fix it! i answered some

lovemelee:

just take your time with it glad to help :)

lovemelee:

oh ok

kacchanforlife:

ive just posted it!

lovemelee:

k